# PerspectiveNet: Multi-View Perception for Dynamic Scene Understanding

### v1

@@ 4 Experiment
Table 1: Evaluation metrics for internal and external scores after the first train.
The same configuration is used for the second training except for the training epoch, which is 4 in the second training. The primary difference between this training and the first one is that we increased the amount of data, which now includes normal multi-view camera data, normal trimmed data, and external data. The result of this training is presented in table Tab. 2. This result shows the significant improvement of the model in both internal and external test datasets. The result has improved by over 2 points for the internal dataset; in the external dataset, this number is just below 6 points. Moreover, after this training, the ability of models to generate descriptions for internal and external data becomes extremely similar. This proved that training models on more data generally leads to better performance.
@@
BLEU 0.2416 0.2363
METEOR 0.4541 0.4598
ROUGE-L 0.4543 0.4576
CIDEr 0.4574 0.3319
S 29.8951 29.6726
Average S 29.7838
Table 2: Evaluation metrics for internal and external scores after the second train.
The training time consuming of each training step is in Tab. 3. Even though half of the dataset used for the first training contains about 4 videos for each sample, this dataset is not very large compared to the full version of the WTS dataset. That explains why the first training is significantly faster than the second one despite we trained the model on this dataset for more epochs than the second one.
Model Training Time (hours)
Model 1 5 hours 15 minutes
Model 2 11 hours 30 minutes
Table 3: Training Time Comparison
The detail of the weight of each module in the model is shown in Tab. 4. Despite the entire model having nearly 3 billion parameters, the visual encoder is not used during training but for the video pre-processing step, which makes the training process only apply to a much smaller model with about 1.5 billion. Moreover, LoRA layers only contain an insignificant number of parameters and the Adapter has roughly 144 million trainable parameters, which is not very large. The weight of the embedding layer in our connector to get information on the object (pedestrian and vehicle) and segment is not included due to the insignificant size.
@@
5 Limitation
Even though this model is small and can generate correct long fine-grain descriptions on multiple-view cameras and videos, this work still has some limitations.
First, this model is constructed to work only with multiple cameras that start at the same time. If any of them start at a different time, the frame feature will not be aligned, which may lead to wrong information comprehension of the model. As a result, the generated descriptions can be unpredictable and the model should not be used for data like that.
Secondly, even though a video is just a list of continuous images, it is not confirmed that training on videos would lead to any improvement in the image description task. Since our model only trained on videos and described a sequence of events without being trained on any image, using this to generate image descriptions is not recommended.
@@ 5 Limitation
6 Conclusion
In this paper, we present PerspectiveNet, a model to generate long and detailed descriptions given videos or multiple-viewpoint videos. The model is a combination of a pretrained vision transformer (ViT), a newly constructed connector from two Perceiver modules, and a large language model. Despite its small size and the language has not been trained on any visual information, this model can still achieve a high score on the WTS dataset and secure a position among the top 5 models currently.

### v2

@@ 4.2 Evaluation Results
Table 1: Evaluation metrics for internal and external scores after the first train (20 epochs in 5 hours).
The same configuration is used for the second training except for the training epoch, which is 4 in the second training. The primary difference between this training and the first one is that we increased the amount of data, which now includes normal multi-view camera data, normal trimmed data, and external data. The result of this training is presented in table 2. This result shows the significant improvement of the model in both internal and external test datasets. The result has improved by over 2 points for the internal dataset; for the external dataset, the training increases the overall score by more than 5 points. Moreover, after this training, the ability of models to generate descriptions for internal and external data becomes extremely similar. This proved that training models on more data generally leads to better performance.
@@
BLEU 0.2663 0.2630
METEOR 0.4637 0.4684
ROUGE-L 0.4587 0.4611
CIDEr 0.9502 1.1400
S 32.0934 32.6614
Average S 32.3774
Table 2: Evaluation metrics for internal and external scores of Our Method (4 more epochs in 11 hours).
The detail of the weight of each module in the model is shown in table 3. Despite the entire model having nearly 3 billion parameters, the visual encoder is not used during training but for the video pre-processing step, which makes the training process only apply to a much smaller model with about 1.5 billion parameters. Moreover, LoRA layers and the Embedding layer only contain an insignificant number of parameters, and the Adapter has roughly 144 million trainable parameters, which is not very large. The embedding layer in our connector, which is used to get information on the object (pedestrian and vehicle) and segment, is not included in this report due to its insignificant number of parameters.
@@
4.3 Methods comparison
Method
BLEU-4
METEOR
ROUGE-L
CIDEr
Score
Model Size
CityLLaVA 0.278 0.477 0.470 1.130 33.43 $\geq$34B
Divide&Conquer 32.89 $\geq$14B
PDVC + CLIP 0.201 0.412 0.442 0.557 29.00 $\geq$2$\times$1B
Multi-perspective + Refinement 22.67 $\geq$2B
Ours 0.2647 0.4660 0.4599 1.0451 32.38 2.7B
Ours (w/o Event Matching) 0.2603 0.4512 0.4461 0.9056 31.20 2.7B
[…]
In this section, we compare our method with some existing models on the WTS dataset. As can be seen from table 4, two models with the best performance used a very large language model to generate descriptions. Meanwhile, methods using smaller LLMs are usually complex, less flexible, and produce worse performance. Our method, despite being unable to overcome large models like CityLLaVA and Devide&Conquer, our model still has significantly better performance than low-resource models and its performance is on par with methods using very large LLMs. Furthermore, unlike methods like PDVC+CLIP Shoman et al. (2024), we do not split our model into two for each scenario, making our method more flexible in general use cases. Lastly, the reported results show that using the Event Matching method increases the model performance by over one point in general, showing its effectiveness and potential application.
5 Conclusion
In this paper, we present PerspectiveNet, a model to generate long and detailed descriptions given videos or multiple-viewpoint videos. The model combines a pretrained vision transformer (ViT), a newly constructed connector from two Perceiver modules, and a large language model. Despite its small size and not being pre-trained on any visual information, this model can still achieve a high score on the WTS dataset, indicating its potential in similar tasks.
Limitation
Even though this model is small and can generate long fine-grain descriptions on multiple-view cameras and videos, this work has some limitations.
First, this model is constructed to work only with multiple cameras that start at the same time. If any of them start at a different time, the frame feature will not be aligned, which may lead to wrong information comprehension by the model. As a result, the generated descriptions can be unpredictable, and the model should not be used for data like that.
@@ 5 Conclusion
Potential risks
This work is LLM-based, which may result in some wrong information, like generating wrong or biased descriptions.